# Characterizing the support of semiclassical measures for higher-dimensional cat maps

### v1

@@ 1.1. Main results
Our first result examines when semiclassical measures must have full support. Suppose $A\in\operatorname{Sp}(2n,\mathbb{Z})$ has a non-unit length eigenvalue. Let $\lambda_{1},\ldots,\lambda_{l}$ be the eigenvalues of $A$ with the largest absolute value. As $A$ is symplectic, $\lambda_{1}^{-1},\ldots,\lambda_{l}^{-1}$ are also eigenvalues of $A$. Let $F_{+}\subset\mathbb{C}^{2n}$ be the sum of generalized eigenspaces corresponding to $\lambda_{1},\ldots,\lambda_{l}$ and let $F_{-}\subset\mathbb{C}^{2n}$ be the sum of generalized eigenspaces corresponding to $\lambda_{1}^{-1},\ldots,\lambda_{l}^{-1}$. Now set $E_{\pm}\coloneqq\{v\in F_{\pm}:\overline{v}=v\}$. For $v\in E_{+}\cup E_{-}$, let $V_{v}\subset\mathbb{Q}^{2n}$ be the smallest rational subspace such that $\mathbb{R}v\subset V_{v}\otimes\mathbb{R}$. Finally, define $\mathbb{T}_{v}\subset\mathbb{T}^{2n}$ to be the subtorus given by the projection of $V_{v}\otimes\mathbb{R}$ onto $\mathbb{T}^{2n}$. Then we have the following characterization of the support of semiclassical measures associated to $A$.
@@
In Appendix B, Anderson and Lemke Oliver show that for 100% of matrices $A\in\operatorname{Sp}(2n,\mathbb{Z})$ (when ordered by any norm on the space of $2n\times 2n$ matrices), the characteristic polynomial of $A^{k}$ is irreducible over the rationals for all $k$. If all powers of $A$ have irreducible characteristic polynomial irreducible over the rationals, then $A|_{E_{\pm}}$ must be diagonalizable over $\mathbb{C}$. We conclude that 100% of $A\in\operatorname{Sp}(2n,\mathbb{Z})$ have semiclassical measures with full support.
@@
We know that $\dim V+\dim V^{\perp\sigma}=n$. Therefore, if $V$ is Lagrangian, $\dim V=2n$ and if $V$ is coisotropic, $n\leq\dim V\leq 2n$. If $V=\mathbb{R}^{2n}$ or if $V$ is a codimension-one subset of $\mathbb{R}^{2n}$, $V$ is automatically coisotropic.
@@ 1.2. Previous results
We briefly review previous results on semiclassical measures. For additional context, see the literature reviews in [dyatlov2021semiclassical]*§1.3 and [Dy]*§3.
@@
Much of the work on semiclassical measures has focused on positive lower bounds for their Kolmogorov–Sinaĭ entropy, known as entropy bounds. In the setting of negatively curved Riemannian manifolds, entropy bounds were established by Anantharaman [anantharaman], Anantharaman and Nonnenmacher [AN], Anantharaman, Koch, and Nonnenmacher [AKN], Rivière [R2, R1], and Anantharaman and Silberman [AS].
@@
Quantum cat maps were introduced in 1980 by Hannay and Berry [HB1980]. Much work has been conducted on arithmetic eigenfunctions, the joint eigenfunctions of a quantum cat map and the Hecke operators. In 2000, Kurlberg and Rudnick showed that the Quantum Unique Ergodicity conjecture holds for arithmetic eigenfunctions in 2 dimensions [KR]. See also the work of Gurevich and Hadani [GH]. In 2010, Kelmer considered higher-dimensional arithmetic eigenfunctions, proving the Quantum Unique Ergodicity conjecture for quantum cat maps that do not have a coisotropic invariant rational subspace [Kel]. As discussed in the previous section, for quantum cat maps that do have an isotropic invariant rational subspace, Kelmer found examples of semiclassical measures supported on proper coisotropic submanifolds of $\mathbb{T}^{2n}$.
@@
Brooks in [Bo10] and Faure and Nonnenmacher in [FN04] proved ‘entropy-like’ bounds for semiclassical measures of 2-dimensional quantum cat maps. In [AKN], Anantharaman, Koch, and Nonnenmacher conjectured that in all dimensions, the optimal lower bound for semiclassical measures of quantum cat maps is half the topological entropy. The entropy of Kelmer’s example [Kel], mentioned above, is half of the topological entropy. Rivière [Ri11] proved a lower bound in higher dimensions that, in some cases, matches the optimal conjecture. Importantly, the Rivière bound matches the entropy for the 2-dimensional example in [FNdB] and our 4-dimensional example in Theorem 1.4. Both examples have half the maximal entropy.
@@
In 2021, Schwartz showed that semiclassical measures for quantum cat maps given by hyperbolic $A\in\operatorname{Sp}(2,\mathbb{Z})$ have full support [Sch]. Soon after, Dyatlov and Jézéquel studied higher-dimensional quantum cat maps given by symplectic matrices $A$ that satisfy a spectral gap condition [dyatlov2021semiclassical]. The authors proved that the support of any semiclassical measure given by such a quantum cat map must contain a subtorus which is dependent on the eigenspace of either the smallest or largest eigenvalue. Furthermore, if the characteristic polynomial of $A$ is irreducible over the rationals, then all semiclassical measures have full support. Both the result of Schwartz [Sch] and of Dyatlov and Jézéquel [dyatlov2021semiclassical] rely on the fractal uncertainty principle of Bourgain and Dyatlov [BD]. However, as we are working in higher dimensions, but do not assume a spectral gap, we cannot use the fractal uncertainty principle of [BD]. To prove Theorem 1.3, we use a more basic uncertainty principle, which easily generalizes to higher dimensions. On the other hand, to prove Theorem 1.1 (1), we generalize on the work of [dyatlov2021semiclassical] using Cohen’s higher-dimensional fractal uncertainty principle [Cohen].
@@ 1.3. Proof outline
We first outline the proofs of Theorems 1.1 and  1.3 (1). Our proof strategy follows [dyatlov2021semiclassical], which in turn was inspired by [DJi][Jin], and [DJN]. Our method diverges from that of [dyatlov2021semiclassical] in several ways, which are noted as they appear in the argument.
Both Theorems 1.1 and 1.3 proceed by contradiction. For $N_{j}\rightarrow\infty$, let $u_{j}\in\mathcal{H}_{N_{j}}(\theta_{j})$ be a normalized sequence of eigenfunctions for $M_{N_{j},\theta_{j}}$ that weakly converges to a semiclassical measure $\mu$. Each space of quantum states $\mathcal{H}_{N_{j}}(\theta_{j})$ is endowed with an inner product $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}_{N_{j}}(\theta_{j})}$, defined in § 2.1. In the proof of Theorem 1.3, we suppose towards a contradiction that $\operatorname{supp}\mu\subset\Sigma$, where $\Sigma$ is the union of finitely many subtori with non-coisotropic tangent space $V$. For Theorem 1.1, we assume that for all $v\in E_{+}\cup E_{-}\setminus\{0\}$ and all $z\in\mathbb{T}^{2n}$, $z+\mathbb{R}v$ intersects $\mathbb{T}^{2n}\setminus\operatorname{supp}\mu$. By Lemma A.1, we know that $\overline{\mathbb{R}v\bmod\mathbb{Z}^{2n}}=\mathbb{T}_{v}$. Thus, by the $A$-invariance of $\mu$, after a contradiction is proven, we can conclude $\overline{\{A^{k}(z+\mathbb{T}_{v}):k\in\mathbb{Z}\}}\subset\operatorname{supp}\mu$.
@@
To reach a contradiction, it suffices to prove $\|B_{\mathcal{Y}}u_{j}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ and $\|B_{\mathcal{X}}u_{j}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ both converge to 0. The support condition on $b_{1}$ gives the decay for $B_{\mathcal{Y}}$. To study $B_{\mathcal{X}}$, we employ the triangle inequality: we bound #$\#\mathcal{X}$ and examine the decay of $\|B_{\mathrm{w}}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ for $\mathrm{w}\in\mathcal{X}$. For the latter, we split $\mathrm{w}$ into two equal parts: $\mathrm{w}=\mathrm{w}_{+}\mathrm{w}_{-}$. Then, we use $\mathrm{w}_{\pm}$ to construct symbols $b_{\pm}$ such the proof of the decay of $\|B_{\mathrm{w}}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ is reduced to the decay of $\|\operatorname{Op}_{h}(b_{+})\operatorname{Op}_{h}(b_{-})\|_{L^{2}\rightarrow L^{2}}$. We show that $\|\operatorname{Op}_{h}(b_{+})\operatorname{Op}_{h}(b_{-})\|_{L^{2}\rightarrow L^{2}}\rightarrow 0$ via different uncertainty principles for Theorems 1.3 and 1.1.
@@
We now summarize the argument for Theorem 1.4. The result generalizes the above-mentioned 2-dimensional counterexample of Faure, Nonnenmacher, and De Bièvre in [FNdB]. Their proof relies on the existence of short periods: for any $A\in\operatorname{Sp}(2,\mathbb{Z})$, there is a sequence $N_{j}\rightarrow\infty$ such that $M_{N_{j},\theta}$ has period $P(N_{j})\sim\frac{2\log N_{j}}{\log\lambda}$. For $N=N_{j}$ and a Gaussian $G_{N}$, they construct a sequence of eigenfunctions by averaging over functions of the form $M_{N,\theta}^{t}G_{N}$ for $|t|\leq P(N)/2$. For $|t|\leq P(N)/4$, $M_{N,\theta}^{t}G_{N}$ weakly converges to $\delta(x,\xi)$, while for $P(N)/4\leq|t|\leq P(N)/2$, $M_{N,\theta}^{t}G_{N}$ weakly converges to $dxd\xi$. Therefore, by averaging, they obtain the semiclassical measure $\tfrac{1}{2}(\delta(x,\xi)+dx_{1}d\xi_{1})$.
@@ 2.1. Semiclassical quantization
[…]
@@
Let $\omega=(y,\eta)\in\mathbb{R}^{2n}$. We call $U_{\omega}\coloneqq\operatorname{Op}_{h}(a_{\omega})$ a quantum translation, where $a_{\omega}(z)\coloneqq\exp(\frac{i}{h}\sigma(\omega,z))$. Noting that $a_{\omega}(z)\in S(1)$, $U_{\omega}$ is well-defined and acts on $\mathscr{S}(\mathbb{R}^{n})$. From [z12semiclassical]*Theorem 4.7,
@@
which depends smoothly on $\theta$. This restriction holds by the definition of $\mathcal{H}_{N}(\theta)$ and the commutation relations given in (2.8).
@@
[…]
@@
By taking $a=1$, we see that every semiclassical measure is a probability measure. From the conjugation condition (2.15), for all $k\in\mathbb{Z}$, $\mu(A^{k}(\Omega))=\mu(\Omega)$. Otherwise put, $\mu$ is $A$-invariant.

### v2

@@ 1.1. Main results
Our first result examines when semiclassical measures must have full support. Suppose $A\in\operatorname{Sp}(2n,\mathbb{Z})$ has a non-unit length eigenvalue. Let $\lambda_{1},\ldots,\lambda_{l}$ be the eigenvalues of $A$ with the largest absolute value. As $A$ is symplectic, $\lambda_{1}^{-1},\ldots,\lambda_{l}^{-1}$ are also eigenvalues of $A$. Let $F_{+}\subset\mathbb{C}^{2n}$ be the sum of generalized eigenspaces corresponding to $\lambda_{1},\ldots,\lambda_{l}$ and let $F_{-}\subset\mathbb{C}^{2n}$ be the sum of generalized eigenspaces corresponding to $\lambda_{1}^{-1},\ldots,\lambda_{l}^{-1}$. Now set $E_{\pm}\coloneqq\{v\in F_{\pm}:\overline{v}=v\}$. For $v\in E_{+}\cup E_{-}$, let $V_{v}\subset\mathbb{Q}^{2n}$ be the smallest rational subspace such that $\mathbb{R}v\subset V_{v}\otimes\mathbb{R}$. Finally, define $\mathbb{T}_{v}\subset\mathbb{T}^{2n}$ to be the subtorus given by the projection of $V_{v}\otimes\mathbb{R}$ onto $\mathbb{T}^{2n}$. In general, we do not have $A\mathbb{T}_{v}=\mathbb{T}_{v}$. Then we have the following characterization of the support of semiclassical measures associated to $A$.
@@
From Lemma B.2, if $n\geq 2$ and the characteristic polynomial of $A$ is irreducible over $\mathbb{Q}$ with the maximal Galois group $S_{2}\wr S_{n}$, then the characteristic polynomial of $A^{k}$ for all $k$ is irreducible. For the definition of the Galois group of a polynomial, see the start of Appendix B.
@@
In Appendix B, Anderson and Lemke Oliver show that for 100% of matrices $A\in\operatorname{Sp}(2n,\mathbb{Z})$ (when ordered by any norm on the space of $2n\times 2n$ matrices), the characteristic polynomial of $A^{k}$ is irreducible over the rationals for all $k$ and $A$ has a non-unit length eigenvalue. If all powers of $A$ have irreducible characteristic polynomial irreducible over the rationals, then $A|_{E_{\pm}}$ must be diagonalizable over $\mathbb{C}$. We conclude that 100% of $A\in\operatorname{Sp}(2n,\mathbb{Z})$ have semiclassical measures with full support.
@@
We know that $\dim V+\dim V^{\perp\sigma}=n$. Therefore, if $V$ is Lagrangian, $\dim V=n$ and if $V$ is coisotropic, $n\leq\dim V\leq 2n$. If $V=\mathbb{R}^{2n}$ or if $V$ is a codimension-one subset of $\mathbb{R}^{2n}$, $V$ is automatically coisotropic.
@@ 1.2. Previous results
We briefly review the literature on semiclassical measures. For additional context, see the literature reviews in [dyatlov2021semiclassical]*§1.3 and [Dy]*§3.
@@
Much of the work on semiclassical measures has focused on positive lower bounds for their Kolmogorov–Sinaĭ entropy, known as entropy bounds. In the setting of negatively curved Riemannian manifolds, entropy bounds were established by Bourgain and Lindenstrauss [BL03], Anantharaman [anantharaman], Anantharaman and Nonnenmacher [AN], Anantharaman, Koch, and Nonnenmacher [AKN], Rivière [R2, R1], and Anantharaman and Silberman [AS]. Using the methods from [anantharaman], it is likely possible that Theorem 1.3 can be improved to show that a semiclassical measure cannot be supported on $\Sigma\times K$, where $\Sigma$ is as in the statement of Theorem 1.3 and $K$ has small topological entropy.
@@
Quantum cat maps were introduced in 1980 by Hannay and Berry [HB1980]. Much work has been conducted on arithmetic eigenfunctions, the joint eigenfunctions of a quantum cat map and the Hecke operators. In 2000, Kurlberg and Rudnick showed that the Quantum Unique Ergodicity conjecture holds for arithmetic eigenfunctions in 2 dimensions [KR]. See also the work of Gurevich and Hadani [GH]. In 2010, Kelmer considered higher-dimensional arithmetic eigenfunctions, proving the Quantum Unique Ergodicity conjecture for quantum cat maps that do not have a coisotropic invariant rational subspace [Kel]. As discussed in the previous section, for quantum cat maps that do have an isotropic invariant rational subspace, Kelmer found sequences of arithmetic eigenfunctions that weakly converge to semiclassical measures supported on proper coisotropic submanifolds of $\mathbb{T}^{2n}$. In 2024, Kurlburg, Ostafe, Rudnick, and Shparlinski showed that in higher dimensions, if there are no coisotropic rational A-invariant subspaces, Quantum Ergodicity holds without the arithmetic assumption  [KORS24].
The benefit of studying arithmetic eigenfunctions is that they satisfy additional symmetries. This idea was further developed by Rivière and Wolf in  [RW25]. They characterized the form of semiclassical measures in the case that their eigenfunctions satisfy extra symmetries (but are not necessarily arithmetic eigenfunctions).
@@
Brooks in [Bo10] and Faure and Nonnenmacher in [FN04] proved ‘entropy-like’ bounds for semiclassical measures of 2-dimensional quantum cat maps. In [AKN], Anantharaman, Koch, and Nonnenmacher conjectured that in all dimensions, the optimal lower bound for semiclassical measures of quantum cat maps is half the topological entropy. The entropy of Kelmer’s example [Kel], mentioned above, is half of the topological entropy. Rivière [Ri11] proved a lower bound in higher dimensions that, in some cases, matches the optimal conjecture. Importantly, the Rivière bound matches the entropy for the 2-dimensional example in [FNdB] and our 4-dimensional example in Theorem 1.4. Both examples have half the maximal entropy. Under additional assumptions on the eigenvalues of the matrix, it is possible to recover Theorem 1.3 from entropy bounds.
@@
In 2021, Schwartz showed that semiclassical measures for quantum cat maps given by hyperbolic $A\in\operatorname{Sp}(2,\mathbb{Z})$ have full support [Sch]. Soon after, Dyatlov and Jézéquel studied higher-dimensional quantum cat maps given by symplectic matrices $A$ that satisfy a spectral gap condition [dyatlov2021semiclassical]. The authors proved that the support of any semiclassical measure given by such a quantum cat map must contain a subtorus which is dependent on the eigenspace of either the smallest or largest eigenvalue. Furthermore, if the characteristic polynomial of $A$ is irreducible over the rationals, then all semiclassical measures have full support. Theorem 1.1 removes the spectral gap condition from the result of  [dyatlov2021semiclassical]. Thus, we characterize the support of semiclassical measures for a larger class of matrices. For example, Theorem 1.1 characterizes the semiclassical measures for matrices of the form $B\oplus B$, where $B\in\operatorname{Sp}(2,\mathbb{Z})$, while  [dyatlov2021semiclassical] does not. Both the result of Schwartz [Sch] and of Dyatlov and Jézéquel [dyatlov2021semiclassical] rely on the fractal uncertainty principle of Bourgain and Dyatlov [BD]. However, as we are working in higher dimensions, but do not assume a spectral gap, we cannot use the fractal uncertainty principle of [BD]. To prove Theorem 1.3, we use a more basic uncertainty principle, which easily generalizes to higher dimensions. On the other hand, to prove Theorem 1.1 (1), we generalize on the work of [dyatlov2021semiclassical] using Cohen’s higher-dimensional fractal uncertainty principle [Cohen].
@@ 1.3. Proof outline
We first outline the proofs of Theorems 1.1 and  1.3 (1). Our proof strategy follows [dyatlov2021semiclassical], which in turn was inspired by [DJi][Jin],  [DJN], and  [anantharaman]. Our method diverges from that of [dyatlov2021semiclassical] in several ways, which are noted as they appear in the argument.
Both Theorems 1.1 and 1.3 proceed by contradiction. For $N_{j}\rightarrow\infty$, let $u_{j}\in\mathcal{H}_{N_{j}}(\theta_{j})$ be a normalized sequence of eigenfunctions for $M_{N_{j},\theta_{j}}$ that weakly converges to a semiclassical measure $\mu$. Each space of quantum states $\mathcal{H}_{N_{j}}(\theta_{j})$ is endowed with an inner product $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}_{N_{j}}(\theta_{j})}$, defined in § 2.1. In the proof of Theorem 1.3, we suppose towards a contradiction that $\operatorname{supp}\mu\subset\Sigma$, where $\Sigma$ is the union of finitely many subtori with non-coisotropic tangent space $V$. For Theorem 1.1, we assume that for all $v\in E_{+}\cup E_{-}\setminus\{0\}$ and all $z\in\mathbb{T}^{2n}$, $z+\mathbb{R}v$ intersects $\mathbb{T}^{2n}\setminus\operatorname{supp}\mu$. These initial assumptions are different from [dyatlov2021semiclassical]. By Lemma A.1, we know that $\overline{\mathbb{R}v\bmod\mathbb{Z}^{2n}}=\mathbb{T}_{v}$. Thus, by the $A$-invariance of $\mu$, after a contradiction is proven, we can conclude $\overline{\{A^{k}(z+\mathbb{T}_{v}):k\in\mathbb{Z}\}}\subset\operatorname{supp}\mu$.
@@
To reach a contradiction, it suffices to prove $\|B_{\mathcal{Y}}u_{j}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ and $\|B_{\mathcal{X}}u_{j}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ both converge to 0. The support condition on $b_{1}$ gives the decay for $B_{\mathcal{Y}}$. To study $B_{\mathcal{X}}$, we employ the triangle inequality: we bound $\mathcal{X}$ and examine the decay of $\|B_{\mathrm{w}}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ for $\mathrm{w}\in\mathcal{X}$. For the latter, we split $\mathrm{w}$ into two equal parts: $\mathrm{w}=\mathrm{w}_{+}\mathrm{w}_{-}$. Then, we use $\mathrm{w}_{\pm}$ to construct symbols $b_{\pm}$ such the proof of the decay of $\|B_{\mathrm{w}}\|_{\mathcal{H}_{N_{j}}(\theta_{j})\rightarrow\mathcal{H}_{N_{j}}(\theta_{j})}$ is reduced to the decay of $\|\operatorname{Op}_{h}(b_{+})\operatorname{Op}_{h}(b_{-})\|_{L^{2}\rightarrow L^{2}}$. We show that $\|\operatorname{Op}_{h}(b_{+})\operatorname{Op}_{h}(b_{-})\|_{L^{2}\rightarrow L^{2}}\rightarrow 0$ via different uncertainty principles for Theorems 1.3 and 1.1, which, in turn, are different from the uncertainty principle used in [dyatlov2021semiclassical].
@@
We now summarize the argument for Theorem 1.4. The result generalizes the above-mentioned 2-dimensional counterexample of Faure, Nonnenmacher, and De Bièvre in [FNdB]. Their proof relies on the existence of short periods: for any hyperbolic $A\in\operatorname{Sp}(2,\mathbb{Z})$, there is a sequence $N_{j}\rightarrow\infty$ such that $M_{N_{j},\theta}$ has period $P(N_{j})\sim\frac{2\log N_{j}}{\log\lambda}$. For $N=N_{j}$ and a Gaussian $G_{N}$, they construct a sequence of eigenfunctions by averaging over functions of the form $M_{N,\theta}^{t}G_{N}$ for $|t|\leq P(N)/2$. For $|t|\leq P(N)/4$, $M_{N,\theta}^{t}G_{N}$ weakly converges to $\delta(x,\xi)$, while for $P(N)/4\leq|t|\leq P(N)/2$, $M_{N,\theta}^{t}G_{N}$ weakly converges to $dxd\xi$. Therefore, by averaging, they obtain the semiclassical measure $\tfrac{1}{2}(\delta(x,\xi)+dx_{1}d\xi_{1})$.
@@ 2.1. Semiclassical quantization
(2.1) $\operatorname{Op}_{h}(a)f(x)\coloneqq\frac{1}{(2\pi h)^{n}}\int_{\mathbb{R}^{2n}}e^{\frac{i}{h}\left\langle x-y,\xi\right\rangle}a\left(\frac{x+y}{2},\xi\right)f(y)dyd\xi,\quad f\in\mathscr{S}(\mathbb{R}^{n}).$
@@
Let $\omega=(y,\eta)\in\mathbb{R}^{2n}$. We call $U_{\omega}\coloneqq\operatorname{Op}_{h}(a_{\omega})$ a quantum translation, where $a_{\omega}(z)\coloneqq\exp(\frac{i}{h}\sigma(\omega,z))$. Noting that $a_{\omega}(z)\in S(1)$, $U_{\omega}$ is well-defined and acts on $\mathscr{S}(\mathbb{R}^{n})$. However, the derivatives of $a_{\omega}$ are not bounded uniformly in $h$. From [z12semiclassical]*Theorem 4.7,
@@
The reduction to $N$ even is for technical simplicity; in the case where $N$ is odd, we pick up the phase constant $e^{-i\pi N\left\langle y,\eta\right\rangle}$.
@@
which depends smoothly on $\theta$ since $\mathcal{H}_{N}(\theta)$ are fibers of the smooth vector bundle $\mathcal{H}_{N}$. This restriction holds by the definition of $\mathcal{H}_{N}(\theta)$ and the commutation relations given in (2.8).
@@
$\left\langle\operatorname{Op}_{N_{j},\theta_{j}}(a)u_{j},u_{j}\right\rangle_{\mathcal{H}_{N_{j}}(\theta_{j})}\rightarrow\int_{\mathbb{T}^{2n}}ad\mu\quad\text{ for all }a\in C^{\infty}(\mathbb{T}^{2n}).$
@@
By taking $a=1$, we see that every semiclassical measure is a probability measure. From the conjugation condition (2.15), for all $k\in\mathbb{Z}$ and Borel sets $\Omega$, $\mu(A^{k}(\Omega))=\mu(\Omega)$. Otherwise put, $\mu$ is $A$-invariant.